2. (4) An airplane is trimmed straight and level at a true airspeed of 50 m/s at sea level. A pull of 5 N is required to holdP = A + BV (9.52)「F-14 ) (9.53)
Answer #1

Giver P=5N V = 70 mis NTR = 5omis Now, P=1/11 % R ) ?) 5^(1-17.)) .:|: -5.2083] AISO, Pe A+ BV? : 5 = A+ B190) .:5 = -5.2083